# The evolution of trust as a cognitive shortcut in repeated interactions

### v2

@@ Abstract
Trust is often thought to increase cooperation. However, game-theoretic models often fail to distinguish between cooperative behaviour and trust. This makes it difficult to measure trust and determine its effect in different social dilemmas. We address this here by formalising trust as a cognitive shortcut in repeated games. This functions by avoiding verifying a partner’s actions once a threshold level of cooperativeness has been observed. We consider trust-based strategies that implement this heuristic, and systematically analyse their evolution across the space of two-player symmetric social dilemma games. We find that where it is costly to verify whether another agent’s actions were cooperative, as is the case in many real-world settings, then trust-based strategies can outcompete standard reciprocal strategies such as Tit-for-Tat in many social dilemmas. Moreover, the presence of trust increases the overall level of cooperation in the population, especially in cases where agents can make unintentional errors in their actions. This occurs even in the presence of strategies designed to build and then exploit trust. Overall, our results demonstrate the individual adaptive benefit to an agent of using a trust heuristic, and provide a formal theory for how trust can promote cooperation in different types of social interaction. We discuss the implications of this for interactions between humans and artificial intelligence agents.
Keywords: Trust modelling, Costly monitoring, Evolutionary game theory, Multi-agent systems, Social dilemma.
@@ 1 Introduction
Trust underlies cooperation in every human society [1, 2], and hence has been called “the glue of healthy societies and the grease of economic productivity” [3]. This implies that the Artificial Intelligence (AI) agents of the future will also need the capability to build trust, in order to cooperate effectively both with each other and with people [4, 5]. But this raises a fundamental question: what exactly is the individual benefit of trusting another agent? By trusting another, an agent exposes itself to risk [6] – the risk that the other agent might not act in an honest and reliable way.
@@
This question has been difficult to answer because there are many different verbal definitions of trust in both social science and artificial intelligence. These definitions can be categorised into: 1. behavioural definitions, which equate trust with more cooperative choices in a game, and 2. psychological definitions, which equate trust with psychological states such as viewing another individual as benevolent and acting with integrity [7]. Prior work on trust in economics [8] and multi-agent systems [9] has typically followed 1., and effectively equated trust with high levels of cooperation. The risky investment decision modelled by the “trust game” [10, 11, 12], which is effectively a one-sided Prisoner’s Dilemma [13], provides a paradigm example. The fundamental problem with this is that it conflates an outcome, cooperation, with one of its causes, trust [14, 15, 16]. As such, these definitions of trust do not allow us to delineate the role of trust in promoting cooperation. Although another (smaller) tradition in multi-agent systems has followed 2., and equipped agents with cognitive mechanisms to make trust decisions [17], this has still not resulted in a general and mechanistic way to measure trust that applies in a wide range of settings. This is because it defines trust relative to an agent’s internal state, which may very well not be observable in an interpretable way, e.g. as is the case with current Large Language Model (LLM) agents, as well as for human agents.
To overcome the problems with both 1. and 2., we need a measure of trust that is behavioural, so that it can be observed regardless of the agent’s architecture (artificial or human), but that does not conflate trust with cooperation. To do so, in this paper we develop an objective measure derived from the adaptive function of trust behaviours, i.e. the individual benefits that trust brings to an agent. We do this by using evolutionary game theory (EGT) to formalise a key theory of trust from social science, originating from the work of Luhmann [6]. Luhmann’s theory defines trust as a heuristic – a decision-making shortcut that people use in social interactions – to avoid having to constantly reason about the likelihood of every possible outcome [6, 7]. When trust is established, the benefit is that people save the time and effort of trying to exhaustively calculate whether or not they should cooperate and take the risk of being exploited.
@@
We turn Luhmann’s theory into a behavioural measure of trust that can apply to both human and artificial agents by building on the theory of repeated games. In repeated games, agents interact with each other several times (rounds). This gives agents the possibility to condition their behaviour on how the other agent has acted previously. For example, by using reciprocal strategies such as “cooperate this time if the other agent cooperated on the previous round”. These reciprocal strategies – where agents can base their knowledge of how the other agent has acted previously on either their own first-hand knowledge (direct reciprocity) or information from third parties (indirect reciprocity) – underpin cooperation in human societies [18]. But crucially, to be able to use reciprocal strategies, an agent needs to know whether the observed actions of another agent are actually cooperative or not. It is traditionally assumed that once an agent observes an action, it can instantly and with zero cost determine whether that observed action was cooperative. However, in reality, verifying whether an observed action was actually cooperative carries an opportunity cost, representing the time and effort spent. For example, in e-commerce it may not be easy for a buyer to tell whether a seller sent them goods of the advertised quality, rather than say a fake, and so choosing to verify this may take a substantial amount of their time. Similarly, verifying the actions of AI agents may involve simulating execution of their source code [19], the cost of which would depend on the power of the simulating agent and the auditing frameworks in place.
@@
Classic game theory has traditionally studied this as a problem of costly monitoring. Some models have considered cases where agents can choose to pay a cost to reveal their co-players’ actions, and have demonstrated folk theorem results about the existence of cooperative equilibria under particular conditions, such as communication [20] or the presence of a random signal [21, 22]. However, these models have focussed on equilibrium outcomes rather than competition amongst different strategies employed by cognitively-bounded agents. Consequently, they have not connected costly monitoring with an individual trust heuristic. By making this connection, we argue that trust can be objectively measured by observing how often an agent chooses not to pay a verification cost, and hence does not invest time and effort in checking whether its fellow agents are cooperating.
@@
We operationalise this functional definition of trust into an agent’s decision-making mechanism by introducing a class of trust-based strategies into repeated games. In contrast to traditional reciprocal strategies such as Tit-for-Tat [23], trust-based strategies only occasionally verify their partner’s actions once a trust threshold has been reached, i.e. after their co-player has made actions that have been verified to be cooperative for a certain number of rounds. By doing so, these strategies reduce the opportunity cost of verifying the action of their partner every round. We previously showed that trust-based strategies evolved in a repeated Prisoner’s Dilemma game, and that their presence increased the frequency of cooperative actions in the overall population of agents. This result held for a wide range of non-zero opportunity costs [24]. But this work left unclear the extent to which this explanation for the evolution of trust would generalise to other common types of social interaction beyond the repeated Prisoner’s Dilemma.
@@
We resolve this here by examining the extent to which trust-based strategies evolve and promote cooperation in a wide range of social interactions, including the full space of symmetric two-player social dilemma games [25, 26], and in settings where agents can make unintentional mistakes in their actions. We find that a trust heuristic is most effective at promoting cooperation in situations where there is a strong temptation to defect. In contrast, trust has the least effect on promoting cooperation in coordination interactions (Stag-Hunt games), although it still evolves in these cases. But in all cases, we find that trust is beneficial in promoting cooperation if there is a moderate probability of agents making unintentional errors in their actions. Overall, our results show that a wide range of repeated social dilemmas select for the evolution of a trust heuristic that avoids always verifying another agent’s actions.
@@ 2 Models and methods
We first describe the model, and then present the methods used to analyse it. We follow the presentation in the Models and methods section of [24], but extend the model to cover different types of game, and introduce the possibility of agents making errors in their actions.
@@ 2.1 Pairwise social dilemmas
We model the interactions between individuals as a pairwise social dilemma, defined by the following parameterised payoff matrix (for row player), as in [25, 26]:
@@
An agent who chooses the cooperate (C) action when their partner chooses the defect (D) action receives the sucker’s payoff $S$, while the agent that defects gains the temptation payoff, $T$. Mutual cooperation (resp., defection) gives a reward $R$ (resp., punishment P) for both agents. This payoff matrix defines the stage game, i.e. the payoffs that result from a single round of the interaction.
Depending on the ordering of the four payoffs, different social dilemmas arise [27, 26]. We use the following scaling of the payoff matrix, where $R=1$, $P=0$, $0\leq T\leq 2$, $1\leq S\leq 1$. Varying $S$ and $T$ capture all possible pairwise social dilemmas. Namely, (see [26])
@@
    Stag Hunt (SH): $1>T>0>S$.
In a repeated game, agents have strategies. These are functions which map from the information that the agent has about the actions its partner has made on previous rounds, to the action that the agent will take in the current round. In a population of $N$ individuals playing a repeated social dilemma as defined above, the fitness of an individual with a strategy A in a population with $k$ As and (Nk)(N-k) Bs is given by
@@
where $\pi_{A,A}(j)$ and $\pi_{A,B}(j)$ denote the payoff obtained by the A individual in a single round, $j$, when it interacts with an individual of strategy A (B, respectively), as defined by the payoff matrix above. The symbol $r$ denotes the total number of rounds of the game. Rather than considering a fixed number of rounds that agents know in advance, we instead make the standard assumption that the game continues for another round with probability $w$. This results in an expected number of $r=(1-w)^{-1}$ rounds per interaction [28]. Throughout this paper, we compute all values of $\Pi$ analytically.
2.2 Conditional strategies and the opportunity cost of verifying a co-player’s actions
In repeated games, agents can condition their strategies on the previous actions of their co-players. Here, we introduce an opportunity cost, denoted by $\epsilon$, which an agent must pay to know whether their partner has just made a cooperative action. The standard Tit-for-Tat (TFT) strategy (“Cooperate on the first round; in subsequent rounds, do the action that the co-player did on the previous round”) pays this cost every round. At the other extreme, fully unconditional strategies, i.e. always cooperate (ALLC) or always defect (ALLD), never pay this cost.
@@ 2.2.1 Trust-based strategies
In previous work, we introduced the concept of trust-based strategies [24]. The TUC (Trust-based Cooperation) strategy starts out playing TFT and verifying the action of its partner every round (paying the cost $\epsilon$ to do so). However, it considers the difference between the number of cooperate versus defect moves its partner has been found to make. Once this exceeds a threshold $\theta$ (known as the trust threshold), then on each subsequent round it cooperates unconditionally with probability 1p1-p (and hence does not pay $\epsilon$). On the other hand, with probability $p$ it verifies the partner’s action on the current round. If as a result the partner is found to defect, it reverts to playing TFT for the remaining rounds (see also Figure 1 of [24]).
@@
For ease of exposition, in the above payoff matrix the payoff for TUC interacting with TUD is written as $\Pi_{TUC,TUD}$, and the payoff for TUD against is written as $\Pi_{TUD,TUC}$. These are defined as follows (as in [24]),
@@
These can be understood as follows. In the first $\theta$ rounds, before the trust threshold is reached, both TUC and TUD play C and keep verifying their partners’ actions. In each of these rounds they therefore both obtain a payoff $R-\epsilon$. Once the trust threshold, $\theta$, is reached, TUC will verify the action of its partner only occasionally with probability $p$, while TUD will switch to defecting. Suppose that in the next round TUC verifies its partner’s action, and so finds it to be defecting. In this case, it obtains $S$ in the current round and $P-\epsilon$ in the remaining rounds, since it reverts to TFT. Otherwise, i.e. if TUC does not verify in that round (with probability 1p1-p), the process above is iterated for the payoffs calculation.
@@ 2.2.2 Error rate
We also allow for the possibility that agents can make unintentional errors in their actions. We assume that with probability $\mu_{\mathrm{e}}$, an agent performs the action opposite to the one given by its strategy (i.e. defects instead of cooperates, or vice versa). Figure 1 illustrates how both TUC and TFT respond to these errors. As the figure illustrates, TUC can be more resilient to these errors if they happen after the trust threshold has been reached, since they will be ignored if they occur on rounds where TUC does not verify.
Figure 1: Trust-based Cooperation (TUC) can be more robust to errors than Tit-for-Tat (TFT). The top panel shows a TUC agent playing against a TFT agent. Both agents start cooperating, and the trust threshold is reached by TUC. After a period of time, TFT makes an error and defects by mistake. However, because TUC has not verified its co-player’s action on this round, this error is overlooked and mutual cooperation resumes for a number of rounds. At a later stage, TFT defects by mistake again. This time, TUC does verify, and reverts to playing TFT for the remaining rounds. By contrast, the lower pane shows that when TFT plays against itself, the first error always causes the breakdown of mutual cooperation.

### v3

# Disentangling trust from cooperation: Trust as reduced monitoring across social dilemmas

@@ Abstract
It is commonly assumed that trust increases cooperation. However, game-theoretic models often fail to distinguish between cooperative actions and trust, making it difficult to independently measure trust and determine how its effects vary in different social dilemmas. To address this, we build on influential theories that equate trust with reduced monitoring of an agent’s actions. We implement this as a heuristic that cognitively bounded agents can use in repeated games to avoid spending time and effort always monitoring their partner. Agents using this heuristic reduce monitoring of a partner’s actions once a threshold level of cooperativeness has been observed – providing a measurable and architecture-agnostic definition of trust. Using evolutionary game theory, we systematically analyse the success of strategies that use this trust heuristic across the entire space of two-player symmetric social dilemma games. We demonstrate that trust-as-reduced-monitoring facilitates cooperation in two different ways. First, when monitoring is costly, trust heuristics allow for higher levels of cooperation in social dilemmas where the temptation to defect is high. Second, when agents can make action errors, trust heuristics promote cooperation even in coordination problems. Our results disentangle trust from cooperation, and provide a behavioural measure of trust that applies across interaction types.
Keywords: Trust, Costly monitoring, Decision heuristic, Evolutionary game theory, Social dilemmas.
@@ 1 Introduction
Trust underlies cooperation in every human society [1, 2], and hence has been called “the glue of healthy societies and the grease of economic productivity” [3]. This implies that the Artificial Intelligence (AI) agents of the future will also need the capability to build trust, in order to cooperate effectively both with each other and with people [4, 5]. However, this raises a fundamental question: what is the individual adaptive benefit of trusting another agent? By trusting another, an agent exposes itself to risk [6] – the risk that the other agent might not act in an honest and reliable way.
@@
This question has been difficult to answer because there are many different verbal definitions of trust in both social science and artificial intelligence. These definitions can be categorised into: 1. behavioural definitions, which equate trust with more cooperative choices in a game, and 2. psychological definitions, which equate trust with psychological states such as viewing another individual as benevolent and acting with integrity [7]. Prior game-theoretic work on trust in economics [8] and multi-agent systems [9] has typically followed 1., and effectively equated trust with high levels of cooperation. The risky investment decision modelled by the “trust game” [10, 11, 12], which is effectively a one-sided Prisoner’s Dilemma [13], provides a paradigm example. The fundamental problem with this is that it conflates an outcome, cooperation, with one of its causes, trust [14, 15, 16]. As such, these definitions of trust do not allow us to delineate the role of trust in promoting cooperation. An alternative tradition in multi-agent systems has followed 2., and equipped agents with cognitive mechanisms to make trust decisions [17]. However, this has still not resulted in a general and mechanistic way to measure trust that applies in a wide range of settings. This is because it defines trust relative to an agent’s internal state, which may not be observable in an interpretable way, e.g. as is the case for current Large Language Model (LLM) agents, as well as for human agents.
To overcome the problems with both 1. and 2., we need a measure of trust that is behavioural, so that it can be observed regardless of an agent’s architecture (artificial or human), but that does not conflate trust with cooperation. In this paper, we develop a behavioural measure derived from the adaptive function of trust behaviours, i.e. the individual benefits that trust brings to an agent. We do this by using evolutionary game theory (EGT) to formalise a key theory of trust – that trusting means reducing monitoring of another agent’s actions. This provides a population-level measure of trust – the frequency of conditional strategies that reduce monitoring. This measure of trust is behavioural, since it only requires knowing whether an agent performs monitoring actions, but does not make reference to an agent’s internal state. Consequently, it is agnostic to agent architecture.
Trust as reduced monitoring fits the intuitive folk psychology rule of “If I trust you then I don’t have to keep checking what you’re doing”. This is in accordance with theories from social science; in particular, Luhmann argues that human agents use a trust heuristic to avoid having to calculate the likelihood of every possible outcome of an interaction [6, 7]. When trust is established, the adaptive benefit is that people save the time and effort of trying to exhaustively calculate whether or not they should cooperate and take the risk of being exploited. This has often been hinted at in the definitions of trust used in artificial intelligence [18]. For example, in Taddeo’s influential definition of e-trust, greater levels of trust entail less effort spent supervising the trusted agent [19]. Similarly, in Castelfranchi and Falcone’s notion of strict trust, trust is antagonistic to monitoring, where monitoring is defined as an action “aimed at ascertaining whether another action has been successfully executed or if a given state of the world has been realized or maintained…” [20, pg. 193]. Trust as a decision to reduce resources (time and effort) invested into monitoring has also been proposed as a pragmatic definition of trust between people and artificial intelligence [21, 22, 23, 24, 25].
Previous work has modelled trust-as-reduced-monitoring using EGT in the context of a repeated Prisoner’s Dilemma game [22]. However, this was restricted in scope to interactions that can be modelled as Prisoner’s Dilemmas, which excludes many common situations where the temptation to defect is less, such as coordination problems. We go beyond this work by providing a game-theoretic formalisation of the trust-as-reduced-monitoring heuristic that applies to many types of interaction settings, while cleanly separating cooperation from trust. To do this, we build on the theory of repeated games.
@@
In repeated games, agents interact with each other several times (rounds). This gives agents the possibility to condition their behaviour on how the other agent has acted previously, for example, by using reciprocal strategies such as “cooperate this time if the other agent cooperated on the previous round”. These reciprocal strategies – where agents can base their knowledge of how the other agent has acted previously on either their own first-hand knowledge (direct reciprocity) or information from third parties (indirect reciprocity) – underpin cooperation in human societies [26]. But crucially, to be able to use reciprocal strategies, an agent needs to know whether the observed actions of another agent are actually cooperative or not. It is traditionally assumed that once an agent observes an action, it can instantly and with zero cost determine whether that observed action was cooperative. In other words, monitoring the actions of a co-player is costless. However, in reality, monitoring whether an observed action was actually cooperative carries an opportunity cost, representing the time, effort and resources spent. For example, in e-commerce it may not be easy for a buyer to tell whether a seller sent them goods of the advertised quality, rather than say a fake, and so choosing to monitor this may take a substantial amount of their time. Similarly, monitoring the actions of AI agents may involve simulating execution of their source code [27], the cost of which would depend on the power of the simulating agent and the auditing frameworks in place. Thus, in most real-world settings, conditional strategies such as Tit-for-Tat would need to pay a monitoring cost in order to know if their co-player cooperated on the previous round or not.
@@
In classic game theory, some models have considered cases where agents can choose whether to pay a cost to reveal their co-players’ actions, and have demonstrated folk theorem results about the existence of cooperative equilibria under particular conditions, such as communication [28] or the presence of a random signal [29, 30]. However, these models have focussed on equilibrium outcomes rather than competition amongst different strategies employed by cognitively bounded agents. Consequently, they have not connected a monitoring cost with an adaptive trust heuristic.
@@
We address this gap by operationalising trust-as-reduced-monitoring into an agent’s decision-making mechanism, through introducing a class of trust-based strategies into repeated games. In contrast to traditional reciprocal strategies that implicitly monitor every round, such as Tit-for-Tat [31], trust-based strategies only occasionally monitor their partner’s actions once a trust threshold has been reached. This trust threshold occurs after they have monitored their co-player for a number of rounds and have found their actions to be cooperative. Trust-based strategies thus reduce the opportunity cost of monitoring their partner’s actions every round.
@@
Our results show the extent to which a trust-as-reduced-monitoring heuristic, as embodied by trust-based strategies, both evolves and promotes cooperation in a wide range of social interactions. We analyse the full space of symmetric two-player social dilemma games [32, 33], and settings where agents can make unintentional mistakes in their actions. We find that trust-as-reduced-monitoring most strongly promotes cooperation in situations where there is a large temptation to defect. In contrast, this trust heuristic has the least effect on promoting cooperation in coordination interactions (Stag-Hunt games), although it still evolves in these cases. But in all cases, including coordination, we find that trust-as-reduced-monitoring is beneficial in promoting cooperation if there is a moderate probability of agents making unintentional errors in their actions. Overall, our results show that a wide range of repeated social dilemmas select for the evolution of a trust heuristic that avoids always monitoring another agent’s actions, and that this promotes cooperation where 1. there is a strong temptation to defect, or 2. agents can make action errors.
@@ 2 Models and methods
We first describe the model and then present the methods used to analyse it. We follow the presentation in the Models and methods section of [22], but extend the model to cover different types of games, and introduce the possibility of agents making errors in their actions.
@@ 2.1 Pairwise social dilemmas
We model the interactions between individuals as a pairwise social dilemma, defined by the following parameterised payoff matrix (for the row player), as in [32, 33]:
@@
An agent who chooses the cooperate (C) action when their partner chooses the defect (D) action receives the sucker’s payoff $S$, while the agent who defects gains the temptation payoff, $T$. Mutual cooperation (resp., defection) gives a reward $R$ (resp., punishment $P$) for both agents. This payoff matrix defines the stage game, i.e. the payoffs that result from a single round of interaction.
Depending on the ordering of the four payoffs, different social dilemmas arise [34, 33]. We use the following scaling of the payoff matrix, where $R=1$, $P=0$, $0\leq T\leq 2$, $1\leq S\leq 1$. Varying $S$ and $T$ captures all possible pairwise social dilemmas. Namely (see [33]):
@@
    Stag-Hunt (SH): $1>T>0>S$.
In a repeated game, agents have strategies. These are functions that map from the information that the agent has about the actions its partner took in previous rounds, to the action that the agent will take in the current round. In a population of $N$ individuals playing a repeated social dilemma as defined above, the fitness of an individual with a strategy A in a population with $k$ As and (Nk)(N-k) Bs is given by
@@
where $\pi_{A,A}(j)$ and $\pi_{A,B}(j)$ denote the payoff obtained by the A individual in a single round, $j$, when it interacts with an individual of strategy A (B, respectively), as defined by the payoff matrix above. The symbol $r$ denotes the total number of rounds in the game. Rather than considering a fixed number of rounds that agents know in advance, we instead make the standard assumption that the game continues for another round with probability $w$. This results in the expected number of $r=(1-w)^{-1}$ rounds per interaction [35]. Throughout this paper, we compute all values of $\Pi$ analytically.
2.2 Conditional strategies and the opportunity cost of monitoring a co-player’s actions
In repeated games, agents can condition their strategies on the previous actions of their co-players. Here, we introduce a monitoring cost, denoted by $\varepsilon$, which an agent must pay to know whether their partner made a cooperative action. The standard Tit-for-Tat (TFT) strategy (“Cooperate on the first round; in subsequent rounds, do the action that the co-player did in the previous round”) pays this cost every round. At the other extreme, fully unconditional strategies, i.e. always cooperate (AllC) or always defect (AllD), never pay this cost.
@@ 2.2.1 Trust-based strategies
Previous work introduced the concept of trust-based strategies for a repeated Prisoner’s Dilemma game [22]. The TUC (Trust-based Cooperation) strategy begins by playing TFT and monitoring the action of its partner every round (paying the cost $\varepsilon$ to do so). However, it considers the difference between the number of cooperate versus defect moves its partner has been found to make. Once this exceeds a threshold $\theta$ (known as the trust threshold), then on each subsequent round it cooperates unconditionally with probability 1p1-p (and hence does not pay $\varepsilon$). On the other hand, with probability $p$ it monitors the partner to determine whether their last action was cooperate or defect. If the partner is found to have defected, it reverts to playing TFT for the remaining rounds (see also Figure 1 of [22]).
@@
For ease of exposition, in the above payoff matrix the payoff for TUC interacting with TUD is written as $\Pi_{TUC,TUD}$, and the payoff for TUD against TUC is written as $\Pi_{TUD,TUC}$. These are defined as follows (as in [22]),
@@
These can be understood as follows. In the first $\theta$ rounds, before the trust threshold is reached, both TUC and TUD play C and keep monitoring their partners’ actions. In each of these rounds they therefore both obtain a payoff $R-\varepsilon$. Once the trust threshold $\theta$ is reached, TUC will monitor the action of its partner only occasionally with probability $p$, while TUD will switch to defecting. Suppose that in the next round TUC verifies its partner’s action and so finds them to be defecting. In this case, it obtains $S$ in the current round and $P-\varepsilon$ in the remaining rounds, since it reverts to TFT. Otherwise, i.e. if TUC does not monitor in that round (with probability 1p1-p), the process above is iterated for the payoffs calculation.
@@ 2.2.2 Error rate
We also allow for the possibility that agents can make unintentional errors in their actions. We assume that with probability $\mu_{\mathrm{e}}$, an agent performs the action opposite to the one given by its strategy (i.e. defects instead of cooperates, or vice versa). Figure 1 illustrates how both TUC and TFT respond to these errors. As the figure illustrates, TUC can be more resilient to these errors if they happen after the trust threshold has been reached, since they will be ignored if they occur on rounds where TUC does not monitor.
Figure 1: Trust-based Cooperation (TUC) can be more robust to errors than Tit-for-Tat (TFT). The top panel shows a TUC agent playing against a TFT agent. Both agents start cooperating, and the trust threshold is reached by TUC. After a period of time, TFT makes an error and defects by mistake. However, because TUC has not monitored its co-player’s action on this round, this error is overlooked and mutual cooperation resumes for a number of rounds. At a later stage, TFT defects by mistake again. This time, TUC does monitor, and reverts to playing TFT for the remaining rounds. By contrast, the lower panel shows that when TFT plays against itself, the first error always causes the breakdown of mutual cooperation.